It is estimated that 40% of households own a riding lawn mower. A sample of 14 households is studied. What is the probability that more than 11 of these own a riding lawn mower?
Answer #1

solution from the given information, we have p= 0.40 q=1-p=0.6 nalu Σ G-P730)=149% 60 ruce Conosco PCY=1)= 149,466.40)%[0.10